OpenStudy (anonymous):

-1 = 5+x over 6 I always have trouble with fraction ones...

OpenStudy (koikkara):

@balooga82 \(\color{blue}{\huge\tt{Welcome~}}\) \(\huge\mathcal{to}\) \(\color{red}{\huge\tt{Open~Study~!}}\) \(-1 = 5+x /6\)

OpenStudy (koikkara):

\(-1 = 5+x /6\) \(-1-5=x/6\) \(-6=x/6\) \(x=?\)

OpenStudy (anonymous):

-36...?

OpenStudy (koikkara):

Yes, you are right... \(\color{blue}{\huge\tt{Nice~}}\) \(\huge\mathcal{to~meet~}\) \(\color{red}{\huge\tt{You~!}}\)

OpenStudy (anonymous):

thank you! do you know how to do this one? v + 9/3 = 8